# SeamPose: Repurposing Seams as Capacitive Sensors in a Shirt for Upper-Body Pose Tracking

### v1

@@ 7.3. Evaluation in Real-World Settings
Moreover, our seam electrodes are capacitive sensors that inherently sense touches and proximity. Because the seam electrodes are very small in size, they are very insensitive to changes in even near proximity (Zhou et al., 2023; Goertz, 2020). However, it remains rather sensitive to touches and direct contact, especially heavy ones causing thread deformation, which may happen during intense exercise or while sleeping. Therefore, to deploy the system for everyday pose-tracking wearing experiences, the system needs to be fine-tuned to be robust to these noises. We plan to take a data-driven approach to collect more data in these scenarios and update our machine-learning model to handle the noise as mentioned above.
@@ 7.4. Different Seam Patterns in Everyday Apparel
SeamPose focuses on leveraging seams in a shirt for upper body motion estimation. Similar to prior work (Zhou et al., 2023), we were only able to evaluate our approach in one shirt with one kind of commonly seen seam placements. However, there are many different clothing patterns. The seam pattern we chose has one of the smallest numbers of seams for long-sleeve garments, so we believe that  SeamPose can be generalized to apparel with more seams, which offer more information on upper-body pose for the model to learn.
As the ablation study shows, the performance decreased as we removed the sleeve electrode but the model can still infer the upper body pose from the seam electrodes around the shoulders. This indicates our approach holds the potential to work on short-sleeve shirts and sleeveless shirts with a slightly worse performance. However, if the shoulder electrodes do not exist, such as on strapped and strapless tops, it is unclear how our system will work. In the future, we plan to integrate sensing seams into more clothing patterns, including tops and bottoms (i.e., pants, shorts, skirts).

### v2

@@ 7.1. Improving Tracking Performance
From prototyping the proof-of-concept shirt, we found the seam signal quality is mainly influenced by (a) the thread selection: conductivity and insulation (Sec. 4.1); and (b) the positions and number of seam electrodes (Sec. 3). We chose to constrain the latter for our design choice of minimal alteration and maximal generalizability and explored the positions and the number of seams with the seam removal ablation study, but an extensive characterization of such factors may improve the information gain.
@@
As shown in Fig. 9, adding more user-specific data to the current user-adaptive model seems to not significantly improve the performance. In the future, we plan to make prototypes of different sizes (i.e., small/medium/large) and introduce user descriptors (e.g., anthropometric data) as inputs to the deep learning pipeline to better learn the variations among different body shapes. Moreover, more data from other users at a significantly larger scale for training the user-independent model may improve both the user-independent and -adaptive performances as shown in many prior systems.
@@ 7.3. Evaluation in Real-World Settings
Moreover, our seam electrodes are capacitive sensors that inherently sense touches and proximity. Because the seam electrodes are very small in size, they are very insensitive to changes in even near proximity (Zhou et al., 2023; Goertz, 2020). At the signal level, there are no changes in the seam signals when the wearer touches electrically grounded and other objects that could cause electromagnetic (EM) interferences, but there are small (relative to body movements and direct electrode contacts) changes when other objects/people approach the electrodes within about 5cm. This finding is consistent with prior work (Zhou et al., 2023). As the tracking level, with our real-time pipeline, there is no obvious performance degradation in the above situations. However, it remains rather sensitive to touches and direct contact, especially heavy ones causing thread deformation, which may happen during intense exercise or while sleeping. Therefore, to deploy the system for everyday pose-tracking wearing experiences, the system needs to be fine-tuned to be robust to these noises. We plan to take a data-driven approach to collect more data in these scenarios and update our machine-learning model to handle the noise as mentioned above.
SeamPose focuses on leveraging seams in a shirt for upper body motion estimation. Similar to prior work (Zhou et al., 2023), we were only able to evaluate our approach in one shirt with one kind of commonly seen seam placements. However, there are many different clothing patterns. For example, seam electrodes around the wrists at the sleeve cuffs could improve/enable forearm/palm orientation tracking, respectively. The seam pattern we chose has one of the smallest numbers of seams for long-sleeve garments, so we believe that  SeamPose can be generalized to apparel with more seams, which offers more information on upper-body pose for the model to learn.
@@ 7.4. Different Seam Patterns in Everyday Apparel
As the ablation study shows, the performance decreased as we removed the sleeve electrode but the model can still infer the upper body pose from the seam electrodes around the shoulders. This indicates our approach holds the potential to work on short-sleeve shirts and sleeveless shirts with a slightly worse performance. However, if the shoulder electrodes do not exist, such as on strapped and strapless tops, it is unclear how our system will work. In the future, we plan to integrate sensing seams into more clothing patterns, including tops and bottoms (i.e., pants, shorts, skirts). Further, we chose the zigzag stitch as a common stitch for seaming on home sewing machines. Zigzag stitches are stretchable and allow natural fabric stretches as the wearer moves. Stretching the fabric with sewn thread in any direction decreases the measured capacitances. After hundreds of sessions with our prototype, there is no performance degradation caused by permanent deformations. From our experience, the stitch type may impact the raw signal qualities but is unlikely to greatly impact the tracking performance. Nevertheless, a formal characterization of stitch types (e.g., overlock stitches common for industrial seaming) remains essential for future work.
@@ 8. Conclusion
Acknowledgements.
This project was partially supported by the National Science Foundation Grant Award No. 2239569. We thank Owen Koonce for appearing in the video figure. We also would like to thank the study participants and the reviewers.